# Enabling stratified sampling in high dimensions via nonlinear dimensionality reduction

### v1

@@ 2.3 A heuristic algorithm to improve the stratification
[…]
However, (M1) and (M2) might be challenging and computationally expensive to solve, so that it might be more effective to simply use a uniform stratification as discussed in Example 2.3. In addition, we would need to solve nonlinear optimization problems in high dimensions, where the variances in the objective functions are not even known explicitly, but must be approximated. Therefore, we propose a heuristic strategy that iteratively refines the strata, while keeping the computational cost limited.
$\mathcal{A}^{(1)}$ =$=\{a_{0}^{(1)},a_{1}^{(1)}\}=\{0,1\},$ (2.32)
$\Lambda^{(1)}$ =$=\{\lambda_{1}^{(1)}\},\quad\lambda_{1}^{(1)}=a_{1}^{(1)}-a_{0}^{(1)}=1,$
$\Sigma^{(1)}$ =$=\{\sigma_{1}^{(1)}\},\quad\sigma_{1}^{(1)}=\operatorname{Var}^{\mu}\left[\mathcal{Q}(X)|\mathcal{F}(\mathcal{E}(X))\in[a_{0}^{(1)},a_{1}^{(1)}]\right]=\operatorname{Var}^{\mu}[\mathcal{Q}(X)].$
$i^{*}=\begin{cases}\operatorname*{arg\,max}_{i\in\{1,\dots,j-1\}}\lambda_{i}^{(j-1)}\sqrt{\sigma_{i}^{(j-1)}},&\text{ for allocation (1)},\\ \operatorname*{arg\,max}_{i\in\{1,\dots,j-1\}}\lambda_{i}^{(j-1)}\sigma_{i}^{(j-1)},&\text{ for allocation (2)},\\ \end{cases}$ (2.33)
@@
$\mathcal{A}^{(j)}$ =$=\{a_{i}^{(j)}\}_{i=0}^{j}=\mathcal{A}^{(j-1)}\cup\{a^{*}\},$ (2.34)
$\Lambda^{(j)}$ =$=\{\lambda_{i}^{(j)}\}_{i=0}^{j},\quad\lambda_{i}^{(j)}=a_{i}^{(j)}-a_{i-1}^{(j)},$
$\Sigma^{(j)}$ =$=\{\sigma_{i}^{(j)}\}_{i=0}^{j},\quad\sigma_{i}^{(j)}=\operatorname{Var}^{\mu}\left[\mathcal{Q}(X)|\mathcal{F}(\mathcal{E}(X))\in[a_{i-1}^{(j)},a_{i}^{(j)}]\right].$
@@ 3 Application to multifidelity estimators
The NeurAM-based stratification presented here can be integrated with other variance reduction methods to achieve even greater precision in the resulting estimates. For example, it can be combined with multifidelity Monte Carlo estimators, as explained next. Let $\mathcal{Q}^{\mathrm{HF}}=\mathcal{Q}$ and assume that a cheap low-fidelity approximation $\mathcal{Q}^{\mathrm{LF}}$ of the high-fidelity model $\mathcal{Q}^{\mathrm{HF}}$ is also known. Let $w=\mathcal{C}^{\mathrm{HF}}/\mathcal{C}^{\mathrm{LF}}$ be the cost ratio between the two model fidelities. The available computational budget $N$, expressed in terms of high-fidelity model evaluations, can then be assembled from the high- and low-fidelity model evaluations $N^{\mathrm{HF}}$ and $N^{\mathrm{LF}}$, respectively, as
@@
It is possible to determine $N^{\mathrm{HF}}$ and $N^{\mathrm{LF}}$ so the resulting multifidelity estimator has minimum variance, or, in other words, an optimal allocation can be computed in closed-form as [23]
@@
where, given the available computational budget in each stratum $\{N_{s}\}_{s=1}^{S}$ such that $N=\sum_{s=1}^{S}N_{s}$ and still following [23], an optimal allocation is given by
@@
NeurAM is applicable to both low- and high-fidelity models to establish a shared space [28]. This space can be used to reparameterize the low-fidelity model, thereby enhancing the correlation $\rho$ between the fidelities. For the remainder of this discussion, we assume that $\mathcal{Q}^{\mathrm{LF}}$ has already undergone reparameterization and demonstrates a strong correlation with $\mathcal{Q}^{\mathrm{HF}}$, satisfying the condition $\rho^{2}>4w/(1+w)^{2}$. Although this reparameterization would provide an additional reduction in variance, it is not needed for the discussion that follows.
@@ 3.1 Allocation strategies and analysis of the variance
Similar to the discussion in Section 2.2, we analyze two allocation strategies for the stratified multifidelity Monte Carlo estimator. An allocation that yields estimators with minimum variance involves selecting $N_{s}$ samples in stratum $s$ according to
@@
×$\hskip 56.9055pt\times\left(\sqrt{1-\left(\sum_{s=1}^{S}\lambda(A_{s})\rho_{s}\right)^{2}}+\sqrt{w\left(\sum_{s=1}^{S}\lambda(A_{s})\rho_{s}\right)^{2}}\right)^{2}$
@@
In Proposition 3.2, the condition for allocation (1) is weaker than the condition for allocation (2), but allocation (2) is easier to apply, since it only requires the probability mass of each strata. Moreover, we notice that Proposition 3.2 only provides sufficient conditions for variance reduction. In particular, a smaller variance can also be obtained even if the assumptions are not satisfied.
The heuristic algorithm proposed in Section 2.3 can still be applied in the context of stratified multifidelity estimators. However, at each iteration, one must remember to always multiply the variance $\widetilde{\sigma}$ of the high-fidelity model in each stratum by a coefficient $\widetilde{\eta}$ given by
@@ 4 Numerical experiments
Figure 2: Contour plot of the model $\mathcal{Q}_{0}$ which is used as a test case for the numerical experiments in Sections 4.1, 4.2, and 4.4.
In this section we analyze the performance and the properties of NeurAM-based stratified estimators. We first study in Section 4.1 how their convergence depends on the quantities $M,K$ and on the number of strata $S$. We remind the reader that $M$ represents the size of the training dataset used to evaluate the loss function (2.7), while $K$ is the number of latent space samples used to approximate the CDF in equation (2.8). Next, in Section 4.2, we show that the heuristic algorithm introduced in Section 2.3 leads to a lower variance with respect to the uniform stratification of the unit interval. In Section 4.3, we compare the proposed stratification strategy with a similar approach, where NeurAM is replaced by active subspace (AS) – a linear dimensionality reduction technique [2]. Then, in Section 4.4, we verify that stratified sampling can be successfully combined with multifidelity estimators for additional variance reduction.
In all the numerical experiments in Sections 4.1, 4.2, 4.3, and 4.4 we consider the domain $\mathbb{D}=[-1,1]^{2}$, the distribution of the input parameters $\mu_{0}=\mathcal{U}([-1,1]^{2})$, and the two-dimensional function $\mathcal{Q}_{0}\colon\mathbb{D}\to\mathbb{R}$
@@
which has already been used in the context of sampling estimators for uncertainty propagation [5, 30, 29, 28]. Its contour plot is shown in Figure 2, and the quantity of interest $q$ can be computed explicitly as
@@
Finally, in Section 4.5 we consider higher-dimensional models and show that our methodology, differently from standard stratified sampling, maintains its desirable properties in high dimensions. The code used to perform the following numerical experiments is available at the GitHub repository https://github.com/AndreaZanoni/StratifiedSamplingNeurAM.
@@
The setup is consistent for all numerical experiments in this section. Specifically, the NeurAM encoder ~$\widetilde{\mathcal{E}}(\cdot;\mathfrak{e})$, decoder $\widetilde{\mathcal{D}}(\cdot;\mathfrak{d})$, and surrogate model $\widetilde{\mathcal{S}}(\cdot;\mathfrak{s})$ are each parameterized as fully connected neural networks with $2$ hidden layers of $8$ neurons each, and $\operatorname{tanh}$ activation. Training is performed by minimizing (2.7) using the Adam optimizer with a fixed learning rate of $10^{-3}$ over 10,000 epochs. Moreover, in each experiment, we compute 1,000 realizations of the stratified Monte Carlo estimator, and use the empirical mean and standard deviation from these realizations to draw a Gaussian approximation in the plots. The computational budget is always denoted by $N$ and does not include the $M$ samples employed to train NeurAM, which could, however, be reused in computationally expensive real applications.
@@ 4.1 Sensitivity to $M$, $K$ and $S$
$M=5$ 1.29e-4 1.00e-4 9.63e-5 1.05e-4
$M=10$ 6.71e-5 7.01e-5 7.07e-5 9.41e-5
$M=50$ 7.96e-5 6.99e-5 6.53e-5 6.38e-5
$M=100$ 7.47e-5 6.85e-5 6.51e-5 6.09e-5
Figure 3: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the proposed estimator $\widehat{q}_{\mathrm{sMC}}$ (solid line), varying the number of data $M=5,10,50,100$ (left), and $K=10^{3},10^{4},10^{5},10^{6}$ (right), used to learn the NeurAM and the CDF, respectively. Top: the gray dash-dotted vertical line represents the exact value of the quantity of interest, while the colored solid vertical lines (left) represent the bias obtained using only the surrogate model $\mathcal{Q}_{\mathrm{S}}$ for the different values of $M$. Bottom: mean squared error of the estimators to be compared with the value 1.93e-4 for Monte Carlo.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the stratified estimators $\widehat{q}_{\mathrm{sMC}}$ (solid line) with both standard and NeurAM-based stratification, varying the number of strata $S=4,9,16,25$. The gray dash-dotted vertical line represents the exact value of the quantity of interest.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: NeurAM-based stratification of the domain for the model $\mathcal{Q}_{0}$, varying the number of strata $S=4,9,16,25$.
We consider the model $\mathcal{Q}_{0}$ in equation (4.1) to study how the performance of the algorithm is affected by the choice of $M$, $K$ and $S$ under a uniform stratification, like in Example 2.3 in Section 2.1. First, we fix the number of strata $S=2$ and the computational budget $N=1024$, and we vary the training dataset size $M=5,10,50,100$, and the number of latent samples $K=10^{3},10^{4},10^{5},10^{6}$, respectively. Approximate distributions and mean squared errors of various estimators are reported in Figure 3, confirming an overall increase in performance with larger training datasets. In particular, the variance of the stratified estimator $\widehat{q}_{\mathrm{sMC}}$ decreases as $M$ increases, until it stabilizes. We remark that the number of samples is not a parameter that can be tuned, but it is usually constrained by the specific application, depending on the available model observations or computational budget.
@@
The hyperparameter $K$ is instead responsible for the bias in the estimation of the quantity of interest, and a smaller $K$ would generally result in a larger bias. However, this is not a limitation of our approach as $K$ can always be chosen sufficiently large. In fact, sampling from the latent space is computationally cheap since, due to Remark 2.2, we can draw samples from $\mu_{0}$ and then apply the encoder $\mathcal{E}$ and the CDF $\mathcal{F}$, which have negligible computational costs. Nevertheless, from the table with the mean squared errors in Figure 3, we notice that increasing $K$ is not always beneficial if $M$ is limited, and this is due to the fact that the one-dimensional manifold is not correctly identified. Yet, the improvement with respect to standard Monte Carlo is still evident, since its mean squared error is 1.93e-4.
We now consider a computational budget $N=3600$, fix $M=100$ and $K=10^{6}$, and vary the number of strata $S=4,9,16,25$, while still adopting a uniform stratification. In Figure 4 we plot the stratified estimator for the different values of $S$, and compare Cartesian and NeurAM-based stratification. The variance of the NeurAM-based estimator is significantly smaller for an increasing number of strata, yielding more precise estimates. Moreover, in Figure 5 we show the stratification produced by NeurAM for an increasing $S$, and observe how each stratum is bounded by the levelsets of $\mathcal{Q}$ in Figure 2. This property is crucial allowing NeurAM-based stratified estimators to scale to high-dimensional problems.

### v2

@@ 2.3 A heuristic algorithm to improve the stratification
$(\mathrm{M}1)$ $\min_{0=a_{0}<a_{1}<\cdots<a_{S-1}<a_{S}=1}\sum_{s=1}^{S}\lambda(A_{s})\sqrt{\operatorname{Var}^{\mu}[\mathcal{Q}(X)|X\in D_{s}]},$ (2.40)
$(\mathrm{M}2)$ $\min_{0=a_{0}<a_{1}<\cdots<a_{S-1}<a_{S}=1}\sum_{s=1}^{S}\lambda(A_{s})\operatorname{Var}^{\mu}[\mathcal{Q}(X)|X\in D_{s}].$
However, (M1) and (M2) might be challenging and computationally expensive to solve, so that it might be more effective to simply use a uniform stratification as discussed in Example˜2.3. In addition, we would need to solve nonlinear optimization problems in high dimensions, where the variances in the objective functions are not even known explicitly, but must be approximated. Therefore, we propose a heuristic strategy that iteratively refines the strata, while keeping the computational cost limited.
$\mathcal{A}^{(1)}$ $=\{a_{0}^{(1)},a_{1}^{(1)}\}=\{0,1\},$ (2.41)
$\Lambda^{(1)}$ $=\{\lambda_{1}^{(1)}\},\quad\lambda_{1}^{(1)}=a_{1}^{(1)}-a_{0}^{(1)}=1,$
$\Sigma^{(1)}$ $=\{\sigma_{1}^{(1)}\},\quad\sigma_{1}^{(1)}=\operatorname{Var}^{\mu}\left[\mathcal{Q}(X)|\mathcal{F}(\mathcal{E}(X))\in[a_{0}^{(1)},a_{1}^{(1)}]\right]=\operatorname{Var}^{\mu}[\mathcal{Q}(X)].$
@@
$\mathcal{A}^{(j)}$ $=\{a_{i}^{(j)}\}_{i=0}^{j}=\mathcal{A}^{(j-1)}\cup\{a^{*}\},$ (2.43)
$\Lambda^{(j)}$ $=\{\lambda_{i}^{(j)}\}_{i=0}^{j},\quad\lambda_{i}^{(j)}=a_{i}^{(j)}-a_{i-1}^{(j)},$
$\Sigma^{(j)}$ $=\{\sigma_{i}^{(j)}\}_{i=0}^{j},\quad\sigma_{i}^{(j)}=\operatorname{Var}^{\mu}\left[\mathcal{Q}(X)|\mathcal{F}(\mathcal{E}(X))\in[a_{i-1}^{(j)},a_{i}^{(j)}]\right].$
@@ 3 Application to multifidelity estimators
The NeurAM-based stratification presented here can be integrated with other variance reduction methods to achieve even greater precision in the resulting estimates. For example, it can be combined with multifidelity Monte Carlo estimators, as explained next. Let $\mathcal{Q}^{\mathrm{HF}}=\mathcal{Q}$ and assume that a cheap low-fidelity approximation $\mathcal{Q}^{\mathrm{LF}}$ of the high-fidelity model $\mathcal{Q}^{\mathrm{HF}}$ is also known. Let $w=\mathcal{C}^{\mathrm{LF}}/\mathcal{C}^{\mathrm{HF}}$ be the cost ratio between the two model fidelities. The available computational budget $N$, expressed in terms of high-fidelity model evaluations, can then be assembled from the high- and low-fidelity model evaluations $N^{\mathrm{HF}}$ and $N^{\mathrm{LF}}$, respectively, as
@@
It is possible to determine $N^{\mathrm{HF}}$ and $N^{\mathrm{LF}}$ so the resulting multifidelity estimator has minimum variance, or, in other words, an optimal allocation can be computed in closed-form as [PWG16]
@@
where, given the available computational budget in each stratum $\{N_{s}\}_{s=1}^{S}$ such that $N=\sum_{s=1}^{S}N_{s}$ and still following [PWG16], an optimal allocation is given by
@@
NeurAM is applicable to both low- and high-fidelity models to establish a shared space [ZGS25]. This space can be used to reparameterize the low-fidelity model, thereby enhancing the correlation $\rho$ between the fidelities. For the remainder of this discussion, we assume that $\mathcal{Q}^{\mathrm{LF}}$ has already undergone reparameterization and demonstrates a strong correlation with $\mathcal{Q}^{\mathrm{HF}}$, satisfying the condition $\rho^{2}>4w/(1+w)^{2}$. Although this reparameterization would provide an additional reduction in variance, it is not needed for the discussion that follows.
@@ 3.1 Allocation strategies and analysis of the variance
Similar to the discussion in Section˜2.2, we analyze two allocation strategies for the stratified multifidelity Monte Carlo estimator. An allocation that yields estimators with minimum variance involves selecting $N_{s}$ samples in stratum $s$ according to
@@
$\hskip 56.9055pt\times\left(\sqrt{1-\left(\sum_{s=1}^{S}\lambda(A_{s})\rho_{s}\right)^{2}}+\sqrt{w\left(\sum_{s=1}^{S}\lambda(A_{s})\rho_{s}\right)^{2}}\right)^{2}$
@@
In Proposition˜3.2, the condition for allocation (1) is weaker than the condition for allocation (2), but allocation (2) is easier to apply, since it only requires the probability mass of each strata. Moreover, we notice that Proposition˜3.2 only provides sufficient conditions for variance reduction. In particular, a smaller variance can also be obtained even if the assumptions are not satisfied.
The heuristic algorithm proposed in Section˜2.3 can still be applied in the context of stratified multifidelity estimators. However, at each iteration, one must remember to always multiply the variance $\widetilde{\sigma}$ of the high-fidelity model in each stratum by a coefficient $\widetilde{\eta}$ given by
@@ 4 Numerical experiments
Figure 2: Contour plot of the model $\mathcal{Q}_{0}$ which is used as a test case for the numerical experiments in Sections˜4.1, 4.2, 4.3, and $4.$
In this section we analyze the performance and the properties of NeurAM-based stratified estimators. We first study in Section˜4.1 how their convergence depends on the quantities $M,K$ and on the number of strata $S$. We remind the reader that $M$ represents the size of the training dataset used to evaluate the loss function (2.7), while $K$ is the number of latent space samples used to approximate the CDF in equation (2.8). Next, in Section˜4.2, we show that the heuristic algorithm introduced in Section˜2.3 leads to a lower variance with respect to the uniform stratification of the unit interval. In Section˜4.3, we compare the proposed stratification strategy with a similar approach, where NeurAM is replaced by active subspace (AS) – a linear dimensionality reduction technique [CDW14]. Then, in Section˜4.4, we verify that stratified sampling can be successfully combined with multifidelity estimators for additional variance reduction.
In all the numerical experiments in Sections˜4.1, 4.2, 4.3, and 4.4 we consider the domain $\mathbb{D}=[-1,1]^{2}$, the distribution of the input parameters $\mu_{0}=\mathcal{U}([-1,1]^{2})$, and the two-dimensional function $\mathcal{Q}_{0}\colon\mathbb{D}\to\mathbb{R}$
@@
which has already been used in the context of sampling estimators for uncertainty propagation [GEG18, ZGS24a, ZGS24b, ZGS25]. Its contour plot is shown in Figure˜2, and the quantity of interest $q$ can be computed explicitly as
@@
Finally, in Sections˜4.5, 4.6, and 4.7 we consider higher-dimensional models, we compare our approach to other sampling-based variance reduction strategies, and study a classical UQ test problem, a Darcy-flow problem, involving partial differential equations, respectively. In particular, we show that our methodology, differently from standard stratified sampling, has the potential to maintains its desirable properties in high dimensions. The code used to perform the following numerical experiments is available at the GitHub repository https://github.com/AndreaZanoni/StratifiedSamplingNeurAM.
@@
The setup is consistent for all numerical experiments in this section. Specifically, the NeurAM encoder $\widetilde{\mathcal{E}}(\cdot;\mathfrak{e})$, decoder $\widetilde{\mathcal{D}}(\cdot;\mathfrak{d})$, and surrogate model $\widetilde{\mathcal{S}}(\cdot;\mathfrak{s})$ are each parameterized as fully connected neural networks with $2$ hidden layers of $8$ neurons each, and $\operatorname{tanh}$ activation. Training is performed by minimizing (2.7) using the Adam optimizer with a fixed learning rate of $10^{-3}$ over 10,000 epochs. Moreover, in each experiment, we compute 1,000 realizations of the stratified Monte Carlo estimator, and use the empirical mean and standard deviation from these realizations to draw a Gaussian approximation in the plots. The computational budget is always denoted by $N$ and does not include the $M$ samples employed to train NeurAM, which could, however, be reused in computationally expensive real applications.
@@ 4.1 Sensitivity to $M$, $K$ and $S$
$M=5$ 1.29e-4 (6.68e-01) 1.00e-4 (5.18e-01) 9.63e-5 (4.99e-01) 1.05e-4 (5.44e-01)
$M=10$ 6.71e-5 (3.48e-01) 7.01e-5 (3.63e-01) 7.07e-5 (3.66e-01) 9.41e-5 (4.88e-01)
$M=50$ 7.96e-5 (4.12e-01) 6.99e-5 (3.62e-01) 6.53e-5 (3.38e-01) 6.38e-5 (3.31e-01)
$M=100$ 7.47e-5 (3.87e-01) 6.85e-5 (3.55e-01) 6.51e-5 (3.37e-01) 6.09e-5 (3.16e-01)
[…]
Figure 3: TOP: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the proposed estimator $\widehat{q}_{\mathrm{sMC}}$ (solid line), varying the number of data $M=5,10,50,100$ (left), and $K=10^{3},10^{4},10^{5},10^{6}$ (right), used to learn the NeurAM and the CDF, respectively. The the gray dash-dotted vertical line represents the exact value of the quantity of interest, while the colored solid vertical lines (left) represent the bias obtained using only the surrogate model $\mathcal{Q}_{\mathrm{S}}$ for the different values of $M$. BOTTOM: Mean squared error (MSE) of the estimators to be compared with the value 1.93e-4 for Monte Carlo obtained with $N=1024$. The numbers in parentheses denote the ratio between the MSE of each estimator and the Monte Carlo reference value.
sMC $S=4$ $S=9$ $S=16$ $S=25$
standard 1.19e-5 (2.09e-01) 1.11e-5 (1.95e-01) 5.55e-6 (9.75e-02) 4.19e-6 (7.36e-02)
NeurAM 6.20e-6 (1.09e-01) 2.08e-6 (3.66e-02) 8.84e-7 (1.55e-02) 4.56e-7 (8.01e-03)
[…]
Figure 4: TOP: NeurAM-based stratification of the domain for the model $\mathcal{Q}_{0}$, varying the number of strata $S=4,9,16,25$. MIDDLE-TOP: Comparison between standard Monte Carlo estimator $\widehat{q}_{\mathrm{MC}}$ (dashed line) and the stratified estimators $\widehat{q}_{\mathrm{sMC}}$ (solid line) with both standard and NeurAM-based stratification. The gray dash-dotted vertical line represents the exact value of the quantity of interest. MIDDLE-BOTTOM: Mean squared error (MSE) of the estimators to be compared with the value 5.69e-5 for Monte Carlo obtained with $N=3600$. The numbers in parentheses denote the ratio between the MSE of each estimator and the Monte Carlo reference value. BOTTOM: MSE as a function of the computational budget $N$.
We consider the model $\mathcal{Q}_{0}$ in equation (4.1) to study how the performance of the algorithm is affected by the choice of $M$, $K$ and $S$ under a uniform stratification, like in Example˜2.3 in Section˜2.1. First, we fix the number of strata $S=2$ and the computational budget $N=1024$, and we vary the training dataset size $M=5,10,50,100$, and the number of latent samples $K=10^{3},10^{4},10^{5},10^{6}$, respectively. Approximate distributions and mean squared errors (MSE) of various estimators are reported in Figure˜3, confirming an overall increase in performance with larger training datasets. In particular, the variance of the stratified estimator $\widehat{q}_{\mathrm{sMC}}$ decreases as $M$ increases, until it stabilizes. We remark that the number of samples is not a parameter that can be tuned, but it is usually constrained by the specific application, depending on the available model observations or computational budget.
@@
The hyperparameter $K$ is instead responsible for the bias in the estimation of the quantity of interest, and a smaller $K$ would generally result in a larger bias. However, this is not a limitation of our approach as $K$ can always be chosen sufficiently large. In fact, sampling from the latent space is computationally cheap since, due to Remark˜2.2, we can draw samples from $\mu_{0}$ and then apply the encoder $\mathcal{E}$ and the CDF $\mathcal{F}$, which have negligible computational costs. Nevertheless, from the table with the mean squared errors in Figure˜3, we notice that increasing $K$ is not always beneficial if $M$ is limited, and this is due to the fact that the one-dimensional manifold is not correctly identified. Yet, the improvement with respect to standard Monte Carlo is still evident, since its MSE is 1.93e-4.
We now consider a computational budget $N=3600$, fix $M=100$ and $K=10^{6}$, and vary the number of strata $S=4,9,16,25$, while still adopting a uniform stratification. In Figure˜4 we plot the stratified estimator for the different values of $S$, and compare Cartesian and NeurAM-based stratification. The variance of the NeurAM-based estimator is significantly smaller for an increasing number of strata, yielding more precise estimates. This behavior can also be seen from the MSE as a function of the computational budget $N=225,450,900,1800,3600$. Moreover, we show the stratification produced by NeurAM for an increasing $S$, and observe how each stratum is bounded by the level sets of $\mathcal{Q}$ in Figure˜2. This property is crucial allowing NeurAM-based stratified estimators to scale to high-dimensional problems.